# Phase locking and multistability in the topological Kuramoto model on cell complexes

### v1

@@ V Class 1: Polygonal faces
The boundary operators are given in Supplementary material. The kernel of the boundary operator ${\boldsymbol{B}}_{[2]}$ is empty, such that we can write ${\boldsymbol{C}}_{[2]}={\boldsymbol{0}}$. The kernel of the boundary operator ${\boldsymbol{B}}_{[1]}^{\top}$ is spanned by the harmonic vector $\boldsymbol{1}$}, such that ${\boldsymbol{C}}_{[1]}^{\top}={\boldsymbol{1}}$. Hence, the set of solutions candidates (16) is given by
@@
Figure 2: Phase locked solutions for Example 1. Filled (empty) markers denote stable (unstable) phase locked solutions. Every solution except $|S_{[0]}^{\bullet}|=|S_{[0]}^{\circ}|=3$ intersects a integer multiple of $2\pi$, i.e. has phase locked solutions. For $|S_{[0]}^{\bullet}|=6,|S_{[0]}^{\circ}|=0$ we find multistability of three phase locked solutions.
Figure 3: Top row: Condition (25) for the ring with 6 edges. Solid lines correspond to the partition $S_{[0]}^{\circ}=\emptyset$, while the dashed to $S_{[0]}^{\circ}=[5,6]$, for selected $\omega_{0}$ values such that ${\boldsymbol{\omega}}=(0,0,0,0,\omega_{0},-\omega_{0})$. Empty (filled) markers correspond to linearly unstable (stable) solutions. As $\omega_{0}$ increases, the number of solutions decreases. The inset shows a curious phenomenon: for a narrow interval of $\omega_{0}$ values, the solution arising from $S_{[0]}^{\circ}=[5,6]$ crosses zero twice, with one solution becoming stable, while $S_{[0]}^{\circ}=\emptyset$ does not intersect zero. Bottom row: The stable solution $S_{[0]}^{\circ}=\emptyset$ (orange line) vanishes at around $\omega_{0}\approx 1.309$, becoming the stable solution $S_{[0]}^{\circ}=[5,6]$. The stable (solid blue line) and unstable solution (dashed blue line) of $S_{[0]}^{\circ}=[5,6]$ vanish in a saddle-node bifurcation around $\omega_{0}\approx 1.327$.
@@
This is a consequence of Eq. (23). The solution with $z^{[-]}=0$ always exists, and other solutions appear in pairs at integer values of $\pm z^{[-]}$, with jumps occurring at $|\frac{s-1}{4}|$, cf. Fig. 4, in line with previous results on the ordinary Kuramoto model [7].
Figure 4: Number of stable solutions $N_{stable}$ depending on ring size $s$ in an all-normal partitioned ring: comparison of numerical results and Eq. (26).
@@ VI Class 2: Platonic solids
For each cell complex, for $n=1$, there are $2^{N_{[2]}}\cdot 2^{N_{[0]}}$ partitions of the sets $S_{[2]}$ and $S_{[0]}$. In larger polyhedra, the analysis of all partitions becomes exhaustive. We therefore limit ourselves in this section to the all-normal partition that provides stable normal solutions, $S_{[2]}^{\bullet}=S_{[2]},S_{[2]}^{\circ}=\emptyset$ and $S_{[0]}^{\bullet}=S_{[0]},S_{[0]}^{\circ}=\emptyset$. Moreover, we once again set ${\boldsymbol{\omega}}={\boldsymbol{0}}$ as our default study case.
We define the cells of the Platonic solids by starting from one initially oriented face, with a fixed vertex order. We then propagate the orientation to neighboring faces by matching shared edges with opposite directions. This way, every face’s orientation becomes consistent with its neighbors. The cells $S_{[0]},S_{[1]},S_{[2]}$, boundary matrices ${\boldsymbol{B}}_{[1]}$, ${\boldsymbol{B}}_{[2]}$ and matrices ${\boldsymbol{C}}_{[1]}$, ${\boldsymbol{C}}_{[2]}$ for each Platonic solid are provided in the Supplementary material. Note that according to our definition above, the boundary matrix from face to volume will always be
${\boldsymbol{B}}_{[3]}={\boldsymbol{1}}\in\mathbb{R}^{N_{[2]}}$
Moreover, the $C$-matrices for $n=1$ are conveniently given by
${\boldsymbol{C}}_{[1]}={\boldsymbol{1}}\in\mathbb{R}^{N_{[0]}},\qquad{\boldsymbol{C}}_{[2]}^{\top}={\boldsymbol{1}}\in\mathbb{R}^{N_{[2]}},$
therefore, in the phase conditions,
${\boldsymbol{1}}^{\top}\,{\boldsymbol{f}}_{[n+1]}({\boldsymbol{\psi}}^{[-]})$ $=2\pi z^{[+]}$
${\boldsymbol{1}}^{\top}\,{\boldsymbol{f}}_{[n-1]}({\boldsymbol{\psi}}^{[+]})$ $=2\pi z^{[-]},$
$z^{[\pm]}$ are scalars.
Moreover, due to symmetry, the elements of the vectors ${\boldsymbol{f}}_{[n+1]}({\boldsymbol{\psi}}^{[-]})$ and ${\boldsymbol{f}}_{[n-1]}({\boldsymbol{\psi}}^{[+]})$ are the same. For ${\boldsymbol{\omega}}=0$, ${\boldsymbol{K}}_{[n]}=\mathrm{diag}(\mathbf{1}_{N_{[0]}})$ and ${\boldsymbol{K}}_{[n+1]}=\mathrm{diag}(\mathbf{1}_{N_{[2]}})$, the set of solution candidates (16) simplifies to
$\psi^{[+]}$ $=\zeta^{[+]}$
$\psi^{[-]}$ $=\zeta^{[-]}$
Since we only consider the all-normal partitions where the $\boldsymbol{f}$}-functions are inverted as $\arcsin$, the phase conditions become
$N_{[2]}\arcsin{\psi^{[+]}}$ $=2\pi z^{[+]}$
$N_{[0]}\arcsin{\psi^{[-]}}$ $=2\pi z^{[-]}.$
Hence, we finally obtain
$\psi^{[+]}$ $=\zeta^{[+]}=\sin\left(\frac{2\pi z^{[+]}}{N_{[2]}}\right),$
$\psi^{[-]}$ $=\zeta^{[-]}=\sin\left(\frac{2\pi z^{[-]}}{N_{[0]}}\right).$
Therefore, the phase-locked states on the edges $n=1$ are determined by the number of nodes $N_{[0]}$ and faces $N_{[2]}$. We can further take into account the constraints on the winding numbers arising from Eq. (23), such that, analogously to a single ring, the lower dimension will contribute with $N_{St,-}$ possible values of $z^{[-]}$ while the upper with $N_{St,+}$ possible values of $z^{[+]}$
$N_{St,-}$ $=1+2\left\lfloor\frac{N_{[0]}-1}{4}\right\rfloor$
$N_{St,+}$ $=1+2\left\lfloor\frac{N_{[2]}-1}{4}\right\rfloor,$
whereby the Cartesian product of the winding numbers gives all possible solutions of $\psi^{[\pm]}$, and the total number of normal fixed points
$N_{stable}=N_{St,-}\times N_{St,+}.$ (27)
@@
The larger the objects, the richer the multistability: the remaining four polyhedra admit respectively 9, 9, 45 and 45 coexisting stable solutions, as shown in Fig. 5. The curious result of coinciding numbers of stable solutions arises from the fact that the cube and octahedron, as well as the dodecahedron and icosahedron, are dual polyhedra: each vertex of one corresponds to a face of the other, and each edge connecting two vertices corresponds to the edge connecting the corresponding faces. Consequently, the product from Eq. (27) remains conserved, whereas the states themselves consist of exchanged positions in $(z^{[-]},z^{[+]})$ or $(\zeta^{[-]},\zeta^{[+]})$, as can be seen from the examples of the phase-locked solutions of the icosahedron and dodecahedron in the $(\zeta^{[+]},\zeta^{[-]})$ plane. Thus, objects that belong to the same symmetry group end up with the same universality pattern.
@@
A detailed graphical exploration of the phase locked solutions is shown in Fig. 5. The cascades of multistability inherited from both lower and higher dimensions become larger for larger objects. On the border of the stability region, we find marginally stable states which do not technically belong to the all-normal partition, since they satisfy $\cos{\boldsymbol{\theta}}^{[-]}=0$.
@@
Figure 5: First row: number of stable phase-locked solutions $N_{\mathrm{stable}}$ for each Platonic solid shown in inset. Second row: phase-locked solutions for all Platonic solids. Same colors and different line styles denote dual Platonic solids to emphasize symmetry. Black dots denote stable phase-locked solutions given by Eq. (22). Other than the tetrahedron, all objects exhibit structural multistability cascades where multistability arises due to interactions from both below and above. In the bottom row, we present these solutions in the $(\zeta^{[+]},\zeta^{[-]})$ plane for the four objects that exhibit multistability.
@@
Having presented the dynamically homogeneous ${\boldsymbol{\omega}}={\boldsymbol{0}}$ case, let us demonstrate how tuning $\boldsymbol{\omega}$} allows us to control the cascade of multistability. We show this phenomenon on the smallest Platonic solid that admits multiple solutions, the cube. For ${\boldsymbol{\omega}}={\boldsymbol{0}}$, the cube has 9 coexisting solutions (cf. Fig. 5). Upon introducing asymmetry to the vector of natural frequencies ${\boldsymbol{\omega}}=(0,0,0,0,0,0,0,0,0,0,+\omega_{0},-\omega_{0})$, the number of stable fixed points decreases gradually, as shown in Fig. 6, disappearing in symmetric pairs or triplets (when a state with the winding number zero is included). In this case, there is no exchange of stability between different partitions as one finds in the ring. States with $S^{\circ}\neq\emptyset$ do not stabilize for any $\omega_{0}$ value.
@@
Figure 6: Reversal of multistability cascade: gradual decrease of number of stable fixed points in the all-normal partition of the cube by introducing asymmetry to the vector of natural frequencies ${\boldsymbol{\omega}}=(0,0,0,0,0,0,0,0,0,0,+\omega_{0},-\omega_{0})$.
@@ VII Class 3: simplexes
Figure 7: Number of stable phase locked states in simplexes up to dimension 5, for each viable $n$ value.
Here, we analyze the stable solutions of the all-normal partitions, $S_{[2]}^{\bullet}=S_{[2]}$ and $S_{[0]}^{\bullet}=S_{[0]}$, for all viable $n$ values for simplexes up to 5 dimensions. For definiteness, we once again set $\omega=0$ and all couplings to unity. Refer to Supplementary material for boundary operators and $\boldsymbol{C}$}-matrices for the 4-simplex and 5-simplex. Our results are summarized in Fig. 7. Note that whenever there is one stable solution, it refers to the trivial solution with winding number(s) zero. We list the winding number vectors for all stable solutions in all considered simplexes.
@@
    The 3-simplex (tetrahedron) consists of 4 vertices, 6 edges and 4 triangular faces, and has been implicitly covered in Sec. VI. It admits single stable states as follows.
@@
VIII Conclusion and Outlook
We presented a new framework for identifying phase-locked states of the topological Kuramoto model on arbitrary cell complexes, the topological nonlinear Kirchhoff conditions algorithm. Having applied our algorithm to elementary motifs, namely polygonal faces, Platonic solids, and simplexes, we demonstrated that cell complexes can generate much richer multistability than simplicial complexes of the same dimension.
Essentially, objects need to be large enough to support multistability of solutions. Solutions are characterized by a set of discrete winding numbers associated to cells of lower and higher dimension. We derive explicit bounds on allowed winding numbers, demonstrating that there needs to be more than 4 cells in the boundary for the system to admit multistability. These bounds also explain the emergence of structural cascades of multistability, where multistability is inherited from lower or upper dimensions, or even both. Furthermore, we find indications of universality classes of multistability, where different objects display identical multistability patterns if their upper and lower boundaries share certain characteristics.
Beyond theoretical interest, these results may be interesting for higher-order oscillator networks in a variety of applications. The cell complexes we consider are elementary building blocks for larger complexes, and our method is applicable to any cell complex, and generalizable to models other than the topological Kuramoto model. The concepts of structural cascades and universality classes in multistability may extend to other dynamical processes on higher-order structures, such as diffusion, consensus, or pattern-forming instabilities. Future work could involve studying complexes comprised of these basic building blocks, both in the context of developing more efficient computational algorithms for larger complexes, and applications to real-world phenomena with non-pairwise interactions. Our findings could be relevant to power-grid stability [2, 51], optical and mechanical oscillator arrays [15, 52], neural ensembles [11], and collective decision-making in multi-agent systems [53], among other applications.

### v2

@@ V Class 1: Polygonal faces
The boundary operators are given in  [topological_repo2025]. We find that the kernel of ${\boldsymbol{B}}_{[2]}$ is empty and the of the boundary operator ${\boldsymbol{B}}_{[1]}^{\top}$ is spanned by the harmonic vector $\boldsymbol{1}$}. Hence, the set of solutions candidates (16) is given by
@@
Figure 3: Phase locked solutions of the topological Kuramoto model for a six-edge ring. Each marker represents a distinct phase-locked solution obtained from the nonlinear Kirchhoff conditions. Filled (empty) markers indicate linearly stable (unstable) solutions. Every partition except $|S_{[0]}^{\bullet}|=|S_{[0]}^{\circ}|=3$ gives rise to phase-locked solutions. For $|S_{[0]}^{\bullet}|=6,|S_{[0]}^{\circ}|=0$, three distinct stable phase-locked solutions coexist, illustrating multistability in the minimal ring motif.
Figure 4: Effect of asymmetry in natural frequencies on stable phase-locked solutions in the ring with six edges. Top row: Plots of condition (25) for two node partitions, $S_{[0]}^{\circ}=\emptyset$ (solid lines) and $S_{[0]}^{\circ}=[5,6]$ (dashed lines) at selected values of $\omega_{0}$ in ${\boldsymbol{\omega}}=(0,0,0,0,\omega_{0},-\omega_{0})$. Filled (empty) symbols mark stable (unstable) solutions. The number of intersections with integer values decreases as $\omega_{0}$ increases, i.e. phase locked states gradually disappear. The inset highlights a narrow interval of $\omega_{0}$ where $S_{[0]}^{\circ}=[5,6]$ crosses zero twice, producing one stable and one unstable solution, while $S_{[0]}^{\circ}=\emptyset$ does not cross it. Bottom row: Dependence of the corresponding phase-locked states on $\omega_{0}$. The stable solution $S_{[0]}^{\circ}=\emptyset$ (orange line) vanishes near $\omega_{0}\approx 1.309$, giving rise to the stable solution $S_{[0]}^{\circ}=[5,6]$ (solid blue line). Stable and unstable (dashed blue line) solutions of $S_{[0]}^{\circ}=[5,6]$ annihilate in a saddle-node bifurcation near $\omega_{0}\approx 1.327$.
@@
This is a consequence of Eq. (23). The solution with $z^{[-]}=0$ always exists, and other solutions appear in pairs at integer values of $\pm z^{[-]}$, with jumps occurring at $|\frac{s-1}{4}|$, cf. Fig. 5, in line with previous results on the ordinary Kuramoto model [manik2017cycle].
Figure 5: Dependence of the number of stable phase-locked states on ring size. The number of stable solutions $N_{stable}$ as a function of ring size $s$ in an all-normal partitioned ring. Symbols show numerical results, and the solid line denotes the analytical prediction from Eq. (26).
@@ VI Class 2: Platonic solids
For each cell complex, for $n=1$, there are $2^{N_{[2]}}\cdot 2^{N_{[0]}}$ partitions of the sets $S_{[2]}$ and $S_{[0]}$. In larger polyhedra, the analysis of all partitions becomes exhaustive. We therefore limit ourselves in this section to the all-normal partition that provides stable normal solutions, $S_{[2]}^{\bullet}=S_{[2]}$ and $S_{[0]}^{\bullet}=S_{[0]}$. Moreover, we once again set ${\boldsymbol{\omega}}={\boldsymbol{0}}$ as our default study case. The boundary matrices are given in [topological_repo2025].
Following algorithm 1 we first parametrize the set of solution candidates. We find that the kernels of the boundary operators are one-dimensional and spanned by the vectors $\boldsymbol{1}$} of appropriate dimension. For ${\boldsymbol{\omega}}=0$, ${\boldsymbol{K}}_{[n]}=\mathrm{diag}(\mathbf{1}_{N_{[0]}})$ and ${\boldsymbol{K}}_{[n+1]}=\mathrm{diag}(\mathbf{1}_{N_{[2]}})$, the set of solution candidates (16) thus simplifies to
${\boldsymbol{\psi}}^{[+]}$ $=\zeta^{[+]}{\boldsymbol{1}}$
${\boldsymbol{\psi}}^{[-]}$ $=\zeta^{[-]}{\boldsymbol{1}}.$
Exploiting symmetry, the phase conditions (22) simplify to
$N_{[2]}\arcsin{\zeta^{[+]}}$ $=2\pi z^{[+]}$
$N_{[0]}\arcsin{\zeta^{[-]}}$ $=2\pi z^{[-]}.$
Hence, we finally obtain the solution
${\boldsymbol{\psi}}^{[+]}$ $=\sin\left(\frac{2\pi z^{[+]}}{N_{[2]}}\right){\boldsymbol{1}},$
${\boldsymbol{\psi}}^{[-]}$ $=\sin\left(\frac{2\pi z^{[-]}}{N_{[0]}}\right){\boldsymbol{1}}.$
We conclude that the normal phase-locked states on the edges $n=1$ are determined by the number of nodes $N_{[0]}$ and faces $N_{[2]}$. We can further take into account the constraints on the winding numbers arising from Eq. (23) and find that $z^{[+]}$ can assume $N_{St,+}=1+2\left\lfloor(N_{[2]}-1)/4\right\rfloor$ different values and $z^{[-]}$ can assume $N_{St,-}=1+2\left\lfloor(N_{[0]}-1)/4\right\rfloor$ different values. All normal fixed points are parameterized by the Cartesian product of the winding numbers $z^{[+]}$ and $z^{[-]}$ such that their total number is given by
$N_{stable}$ $=N_{St,-}\,N_{St,+}$ (27)
$=\left(1+2\left\lfloor\frac{N_{[0]}-1}{4}\right\rfloor\right)\left(1+2\left\lfloor\frac{N_{[2]}-1}{4}\right\rfloor\right).$
@@
The larger the objects, the richer the multistability: the remaining four polyhedra admit respectively 9, 9, 45 and 45 coexisting stable solutions, as shown in Fig. 6. The curious result of coinciding numbers of stable solutions arises from the fact that the cube and octahedron, as well as the dodecahedron and icosahedron, are dual polyhedra: each vertex of one corresponds to a face of the other, and each edge connecting two vertices corresponds to the edge connecting the corresponding faces. Consequently, the product from Eq. (27) remains conserved, whereas the states themselves consist of exchanged positions in $(z^{[-]},z^{[+]})$. Thus, objects that belong to the same symmetry group end up with the same universality pattern.
@@
A detailed graphical exploration of the phase-locked solutions is shown in Fig. 6. The cascades of multistability inherited from both lower and higher dimensions become larger for larger objects. On the border of the stability region, we find marginally stable states which do not technically belong to the all-normal partition, since they satisfy $\cos{\boldsymbol{\theta}}^{[-]}=0$.
@@
Figure 6: Multistability in Platonic solids (edge dynamics, $n=1$). Top row: number of stable phase-locked states $N_{\mathrm{stable}}$ for each Platonic solid (all-normal partition, ${\boldsymbol{\omega}}=0$). Middle row: phase-locked solutions as solutions of the topological nonlinear Kirchhoff conditions Eq. (22) (black dots denote stable solutions). Other than the tetrahedron, all objects display structural cascades of multistability where stable states are inherited from both lower and higher dimensions. Dual Platonic solids are shown in matching colors with different line styles to emphasize their symmetry. Bottom row: stable phase-locked solutions in the $(\zeta^{[+]},\zeta^{[-]})$ plane for the four solids that exhibit multistability.
@@
Having presented the dynamically homogeneous ${\boldsymbol{\omega}}={\boldsymbol{0}}$ case, let us demonstrate how tuning $\boldsymbol{\omega}$} allows us to control the cascade of multistability. We show this phenomenon on the smallest Platonic solid that admits multiple solutions, the cube. For ${\boldsymbol{\omega}}={\boldsymbol{0}}$, the cube has 9 coexisting solutions (cf. Fig. 6). Upon introducing asymmetry to the vector of natural frequencies ${\boldsymbol{\omega}}=(0,0,0,0,0,0,0,0,+\omega_{0},0,0,+\omega_{0})$, the number of stable fixed points decreases gradually, as shown in Fig. 7, disappearing in symmetric pairs or triplets (when a state with the winding number zero is included). In this case, there is no exchange of stability between different partitions as one finds in the ring. States with $S^{\circ}\neq\emptyset$ do not stabilize for any $\omega_{0}$ value.
@@
Figure 7: Reversal of the multistability cascade in the cube. Gradual reduction in the number of stable fixed points for the all-normal partition of the cube as asymmetry is introduced into the vector of natural frequencies, ${\boldsymbol{\omega}}=(0,0,0,0,0,0,0,0,+\omega_{0},0,0,+\omega_{0})$. The plot shows how increasing $\omega_{0}$ reverses the structural cascade of multistability observed in the symmetric case.
@@ VII Class 3: simplexes
Figure 8: Number of stable phase-locked states in simplexes up to dimension 5. For each cell dimension $n\in\{0,...,d-1\}$, the number of stable (normal) phase-locked states obtained from the topological nonlinear Kirchhoff conditions with ${\boldsymbol{\omega}}=0$ is shown. Multistability occurs only when at least one relevant boundary contains more than four cells; the 4-simplex is the smallest case satisfying this criterion. Pairs of $n$ and $d$ that exhibit multistability are highlighted in grey.
Here, we analyze the stable solutions of the all-normal partitions, $S_{[2]}^{\bullet}=S_{[2]}$ and $S_{[0]}^{\bullet}=S_{[0]}$, for all viable $n$ values for simplexes up to 5 dimensions. For definiteness, we once again set ${\boldsymbol{\omega}}=0$ and all couplings to unity. The explicit form of the boundary operators and the kernel matrices are given in  [topological_repo2025]. Our results are summarized in Fig. 8. Note that whenever there is one stable solution, it refers to the trivial solution with winding number(s) zero. We list the winding number vectors for all stable solutions in all considered simplexes.
@@
    The 3-simplex (tetrahedron) consists of 4 vertices, 6 edges and 4 triangular faces. It admits single stable states as follows.
@@
We can generalize our result on the dimension of $\boldsymbol{z}$$^{[\pm]}$ to simplexes of arbitrary dimension $d$ and obtain
$\dim{\boldsymbol{z}}^{[+]}$ $=\dim\ker({\boldsymbol{B}}_{n+1})=\binom{d}{n+2}$
$\dim{\boldsymbol{z}}^{[-]}$ $=\dim\ker({\boldsymbol{B}}_{n}^{\top})=\binom{d}{n-1}$ (28)
with $\dim{\boldsymbol{z}}^{[-]}=0$ for $n=0$ (see appendix for a proof). The elements of $\boldsymbol{z}$$^{[\pm]}$ are integers. One can then apply the equations (23) to compute upper bounds for the number of allowed winding numbers. Deciding which vectors $\boldsymbol{z}$$^{[\pm]}$ are achievable is assisted by symmetry arguments: simplices with ${\boldsymbol{\omega}}={\boldsymbol{0}}$ and equal coupling are symmetric to vertex permutations. Hence, if a certain $\boldsymbol{z}$$^{[\pm]}$ is achievable, then any permutation of its components, should also be achievable.
VIII Discussion
Our results establish a general theoretical framework for studying multistability in oscillator networks with higher-order interactions, integrating concepts from network science, geometry, and nonlinear dynamics. The topological nonlinear Kirchhoff conditions algorithm allows one to identify all phase-locked states of the topological Kuramoto model on arbitrary cell complexes. By applying it to minimal motifs (polygonal rings, Platonic solids, and simplexes), we showed that multistability emerges when boundaries contain more than four cells. Structural cascades of multistability then arise, with stable states inherited from lower and/or upper dimensions. Moreover, dual Platonic solids display identical multistability patterns, suggesting the existence of universality classes determined by the boundary structure.
Future research could explore how these structural cascades manifest in heterogeneous systems, and how combining the elementary motifs studied here gives rise to multistability in larger complexes. Beyond theoretical interest, these findings are relevant for real-world systems where higher-order couplings are intrinsic, such as neuronal networks  [Breakspear2010], optical and mechanical oscillator arrays  [Nixon2013, Zhang2012], power-grid stability  [dorfler2014synchronization, Motter2013], and collective decision-making in multi-agent systems [OlfatiSaber2007]. The framework may be applicable to other dynamical processes on cell complexes, including diffusion, consensus, and pattern-forming instabilities.